OpenStudy (anonymous):

Write an equation in slope intercept form of the line passing through (4,1) and (5-1)

OpenStudy (ivettef365):

first you need to find the slope using formula y2 - y1 ------- x2 - x1

OpenStudy (anonymous):

i knew you would put that ha

OpenStudy (ivettef365):

I see you got me

OpenStudy (anonymous):

ok is it y=-4x+1?

OpenStudy (anonymous):

....

OpenStudy (ivettef365):

lets find the slope 1 2 x,y x,y you have points (4,1) and (5,-1) -1-1 -2 ----- = ----- = -2 5-4 1

OpenStudy (ivettef365):

so slope is -2, now to find the equation of the line, you need in form y = mx + b where m= slope b = y-int so all we need is to find b we get any point, lets take (4,1) 1 = -2(4) + b 1 = -8 + b 1+8 = b b = 9

OpenStudy (anonymous):

y=?-2x+9

OpenStudy (ivettef365):

correct y = -2x +9

OpenStudy (anonymous):

ok mind helping with like 1 or 2 more?

OpenStudy (anonymous):

not the same type of questions though